# A two-body femtoscopy approach to the proton-deuteron correlation function

### v1

@@ V Note added
After the completion of this work and during the writing of the manuscript, the preprint [51] appeared. The motivation and the methodology of this papers coincides with the ones here, and the results are entirely compatible with ours. We fully agree with the conclusions of [51] noting that the Lednický-Lyuboshitz approximation might not be reliable for small sources and, in particular, for correlation functions involving the deuteron.
The potential used in [51] is a combination of square wells, while here we have used Woods-Saxon and a Gaussian profiles. For comparison, we present in Fig. 5 the $pd$ correlation function with a source radius of $r_{0}=3$ fm taken from [51] together with our calculations with the WS and Gaussian potentials. All calculations are taken up to $L=2$. We observe small differences tied to the particular form of the potentials that provide a different behavior of the wave functions for $r\lesssim 3$ fm.
Figure 5: Comparison of the $pd$ correlation function from the solution of the 2-body Schrödinger equation with different potentials, the Woods-Saxon potential from [41], the Gaussian potential parametrized in this work, and a combination of squares wells from the recent Ref. [51].

### v2

@@ V Note added
After the completion of this work and during the writing of the manuscript, the preprint [54] appeared. The motivation and the methodology of this paper coincides with the one used here, and the results are entirely compatible with ours. We fully agree with the conclusions of Ref. [54] that the Lednický-Lyuboshitz approximation might not be reliable for small sources and, in particular, for correlation functions involving the deuteron.
The potential used in Ref. [54] is a combination of square wells, while here we have used Woods-Saxon and Gaussian profiles. For comparison, we present in Fig. 5 the $pd$ correlation function with a source radius of $r_{0}=3$ fm taken from Ref. [54] together with our calculations with the WS and Gaussian potentials. All calculations are taken up to $L=2$. We observe small differences due to the particular form of the potentials that provide a different behavior of the wave functions for $r\lesssim 3$ fm.
Figure 5: Comparison of the $pd$ correlation function from the solution of the 2-body Schrödinger equation with different potentials, the Woods-Saxon potential from Ref. [42], the Gaussian potential parametrized in this work, and a combination of squares wells from the recent Ref. [54].
@@ Acknowledgements.
Appendix A Lednický-Lyuboshitz approximation
In the Lednický-Lyuboshitz approximation (LL) [55], the pair wave function $\psi(\bm{r};\bm{k})$ going into the Koonin-Pratt formula is taken in its asymptotic form, even for small values of $\bm{r}$. In the nonrelativistic case and without the electromagnetic interaction, this amounts to using [46],
$\psi(\bm{r};\bm{k})=e^{i\bm{k}\cdot\bm{z}}+f(k)\frac{e^{ikr}}{r}\ ,$ (13)
where $f(k)\equiv f(\bm{k}\rightarrow\bm{k^{\prime}})|_{|\bm{k}|=|\bm{k}^{\prime}|}$ is the scattering amplitude. This quantity contains all the information on the strong interaction, and in the standard use of the LL approximation, only the $L=0$ partial wave is taken. See, for example, the application to the proton-deuteron case by the ALICE collaboration in Ref. [28]. In this particular system, the Coulomb interaction should also be considered and the total wave function $\Phi(r,z;k)$ needs to be modified according to the Coulomb case [46, 53],
$\Phi(r,z;k)=\Phi^{\textrm{C}}(r,z;k)+e^{-\pi\gamma/2}\Gamma(1+i\gamma)G(\gamma)^{1/2}f_{C}(k)\frac{iF_{0}(\gamma;kr)-G_{0}(\gamma;kr)}{r}\ ,$ (14)
where $\Phi^{C}(r,z;k)$ is the total Coulomb wave function given in Eq. (9), $\gamma$ is the Sommerfeld factor $\gamma=Z_{1}Z_{2}\mu\alpha/k$, $G(\gamma)$ is the Gamow factor,
$G(\gamma)=\frac{2\pi\gamma}{e^{2\pi\gamma}-1}\ ,$ (15)
$F_{0}$ and $G_{0}$ are the regular and irregular $L=0$ Coulomb wave functions, respectively, and $f_{C}(k)$ is the scattering amplitude in the presence of the electromagnetic force. In Ref. [55] the low-energy expansion of $f_{\rm{C}}(k)$ is also used,
$f_{\rm{C}}^{-1}(k)=-\frac{1}{a_{0}}+\frac{1}{2}d_{0}k^{2}-2k\gamma h(\gamma^{-1})-ikG(\gamma)\ ,$ (16)
where
$h(x^{-1})=-\log(|x|)+\frac{1}{2}\psi(1-ix)+\frac{1}{2}\psi(1+ix)\ ,$ (17)
with $\psi(x)$ the digamma function.
Because the presence of the irregular function $G_{0}(\gamma,kr)$ in Eq. (14) forbids the extension of the asymptotic form to small values of $kr$, in Ref. [55] a further approximation is made for $\gamma kr\ll 1$ in which ${{}_{1}}F_{1}(-i\gamma;1;ik(r-z))\simeq 1$, $F_{0}(\gamma;kr)\simeq G(\gamma)^{1/2}\sin(kr)$ and $G_{0}(\gamma;kr)\simeq-G(\gamma)^{-1/2}\cos(kr)$.
With these modifications, and using the Gaussian source in Eq. (7) we have obtained an analytic formula for the LL approximation with Coulomb interaction in terms of the strong scattering length and scattering range,
$C_{\rm{LL}}(k)=G(\gamma)\left\{1+\frac{2\textrm{Re}f_{\rm{C}}}{\sqrt{\pi}r_{0}}F_{1}(2qr_{0})-\textrm{Im}f_{\textrm{C}}\frac{G(\gamma)}{r_{0}}F_{2}(2qr_{0})+\frac{|f_{C}|^{2}}{2r_{0}^{2}}\left[1+\frac{1}{2}[G(\gamma)^{2}-1](1-e^{-4q^{2}r_{0}^{2}})\right]\right\}\ ,$ (18)
[…]
To apply Eq. (18) to our results, we first extract the low-energy parameters ($a_{0}^{1/2},d_{0}^{1/2},a_{0}^{3/2},d_{0}^{3/2}$) by adjusting the following expression
$k\cot\delta^{J}_{0}(k)=-\frac{1}{a^{J}_{0}}+\frac{1}{2}d_{0}^{J}k^{2}-2q\gamma h(\gamma^{-1})\ ,$ (19)
to the phase shifts shown in Fig. 1, namely the original Arvieux data [33] and those obtained using the Woods-Saxon and Gaussian potentials (which we recall were precisely fitted to the former data). For further reference, we note that we use the same sign convention for the scattering length as in Ref. [33], consistent also to the sign in the expression of Eq. (16) (which is the opposite convention as in our Ref. [53]).
We summarize the resulting low-energy parameters in Table 3.
$L=0,S=1/2$ $L=0,S=3/2$
$a_{0}$ (fm) $d_{0}$ (fm) $a_{0}$ (fm) $d_{0}$ (fm)
Arvieux $(2.72\pm 0.08)$ $(2.31\pm 0.10)$ $(11.93\pm 0.25)$ $(2.50\pm 0.04)$
Woods-Saxon $3.68$ $1.44$ $11.94$ $2.47$
Gaussian $4.07$ $1.71$ $10.95$ $2.52$
Table 3: Extracted low-energy parameters in the data and models used in this work by fitting the $pd$ phase shifts according to Eq. (19). “Arvieux” refers to the experimental data given in Ref. [33], and “Woods-Saxon” and “Gaussian” to the phase shifts obtained with these two potentials.
The straight fit to Arvieux’s data—within the same momentum range as indicated in Ref. [33]—is fully consistent with the numbers quoted in that reference. The fits to the phase shifts generated by the Woods-Saxon and Gaussian potentials are useful to quantify the systematic uncertainty due to the choice of the microscopic model. We notice that the stronger dispersion of the extracted low-energy parameters appears in the $S=1/2$ channel, while those for $S=3/2$ are pretty consistent among each other.
We finally implement these low-energy scattering data into the LL formula in Eq. (18) and generate the $pd$ correlation function in this approximation. We present our results in Fig. 6.
Figure 6: Proton-deuteron correlation function in the LL approximation using the low-energy parameters of the data and models shown in Table 3. See details in the main text.
[…]